OpenStudy (anonymous):

help me again any1 : \[\color{red}{\text{b^2+5>9b+12}}\]

OpenStudy (anonymous):

\[b^2+5>9\] this is an inequality .this is satisfied if : 1) \[b>9, b<1 \] 2) \[b=10,b=-1\] 3) \[b>8 , b<0\]

OpenStudy (anonymous):

b^2-9b+5-12=0 b^2-9b-7=0 -----> b1=(9/2)+4sqrt(2) , b2=(9/2)-4sqrt(2) so the solution area would be as (-infinity,b2) U ( b1,+inifinity)

OpenStudy (anonymous):

didnt understand raheen, what is the answer finally

OpenStudy (anonymous):

\[(-\infty,(9/2)-4\sqrt{2}] U [(9/2)+4\sqrt{2},+\infty)\]